OpenStudy (anonymous):

f(x)=x/2-x. Find f(x+2)

OpenStudy (cwrw238):

replace x in f(x) by x+2

OpenStudy (cwrw238):

= (x+2)/ [2 - (x+2)] = (x+2) / -x = -(x+2) / x

OpenStudy (phi):

cwrw238 is for \[ f(x)= \frac{x}{2-x} \] (which should be typed in as f9x)= x/(2-x) ) if we interpret the question as typed: \[ f(x)= \frac{x}{2} -x \] we would (obviously) get a different answer. But the idea is the same: replace x with (x+2)

OpenStudy (anonymous):

= (x+2)/ [2 - (x+2)] = (x+2) / -x how did that happen?

OpenStudy (anonymous):

as in what happened to the 2s in the denominator???

OpenStudy (phi):

2-(x+2) is the same as 2 + -1*(x+2) distribute the -1 to get 2+ -1*x + -1*2 or 2-x-2 2-x-2 is the same as 2 + (-x) + (-2) and when adding, order can be changed: 2 + (-2) + (-x) = 0+ (-x) = -x

OpenStudy (anonymous):

sweet.

OpenStudy (phi):

your are a connoisseur!